Sammamish Brick, Inc., manufactures bricks using clay deposits on the companys property. Raw clays are blended and then extr1b. Is there sufficient capacity in the kiln to satisfy demand for all products? O Yes O No 2a. Calculate the profitability i2b. What is the production plan for the year that would maximize the companys profit? (Round answer to the nearest complete3. What would be the total contribution margin for the production plan you have proposed? Traditional Brick Textured Facing C4. The kiln could be operated for more than 3,325 hours per year by running it after normal working hours. Up to how much per6. Salespersons are currently paid a commission of 5% of gross revenues. Will this motivate the salespersons to make the righ
Answer #1

1a.

Traditional Textured Cinder Roman Total
Brick Facing Block Brick
Annual demand in pallets 125 145 135 185
Hours required in the drying kiln per pallet 10 10 4 5
Total hours required in the drying kiln 1250 1450 540 925 4165

1b. No

There isn't sufficient capacity in the kiln since 4165 hours are required in the kiln to satisfy the demand while only 3325 hours are available.

2a.

Traditional Textured Cinder Roman
Brick Facing Block Brick
Contribution margin per pallet 590 700 480 570
Hours required in the drying kiln per pallet 10 10 4 5
Profitability index 59 70 120 114

2b.

Cinder Block 135 pallets
Roman Brick 185 pallets
Textured Facing 145 pallets
Traditional Brick* 41 pallets

*Based on the profitability index, the products will be produced starting from product with the highest profitability index progressing to the ones with the lower profitability index. Thus, Cinder Block, followed by Roman Brick, and Textured Facing will be produced first and the remaining drying kiln hours available will then be utilized for manufacturing Traditional Brick which has the lowest profitability index.

Number of pallets of Traditional Brick produced = Remaining hours/Hours required in drying kiln per pallet = [3325 - (540 + 925 + 1450)]/10 = 410/10 = 41 pallets.

Per HOMEWORKLIB RULES, the first 4 sub-parts have been answered. Please post the remaining separately. Thank you.